# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ 5.3. Automorphic description of Eisenstein cohomology
In order to describe the main result of this section, we need to consider a modification of the intertwining operators $T(\phi,s)$ introduced in the previous subsection.
@@
Let $\lambda$ be a character of $\mathrm{T}^{\prime}$ and $w$ an element of $W$ of length $\geq 2$. Let $\phi$ be an Hecke character of $F$ of type $w\star\lambda$. Let $s\in{\mathbb{C}}$ belong to the half-plane of convergence of the Euler product of $L(\phi,s)$. For every place $v$, we define
@@
    If $s$ satisfies the hypotheses, i.e. $\mbox{Re}\ s$ satisfies the bounds of Lemma 5.4.(1), then both the operator $T(\phi_{v},s)$ and the quantity $c_{v}(\phi,s)$ are defined and non-zero (cfr. [17, Lemma 2.3.1]).
@@
For the rest of this section, we suppose that $\lambda$, $w$ and $\phi$ are as in the definition above.
@@
    For each $\phi\in\Phi_{\mbox{\tiny{np}}}$, we denote by $\mathop{\rm Eis}\nolimits_{\phi}$ the operator induced in $(\mathfrak{g},\tilde{K}_{\infty})$-cohomology by the the operator $\mathop{\rm Eis}\nolimits_{\phi}$ of Definition 5.9.(1).
  2. (2)
    For each $\phi\in\Phi_{\mbox{\tiny{p}}}$, we denote by $\mathop{\rm Eis}\nolimits^{\prime}_{\phi}$ the restriction to $I^{\prime}_{\phi}$ of the operator $\mathop{\rm Eis}\nolimits_{\phi}$ induced in $(\mathfrak{g},\tilde{K}_{\infty})$-cohomology by the operator $\mathop{\rm Eis}\nolimits_{\phi}$ of Definition 5.9.(2).
Definition 5.11.
  1. (1)
    We denote $I^{\prime}_{\phi}$ the submodule of $I_{\phi}$ generated by those tensors, whose component at $v$ belongs to the kernel of $T^{\mathop{\rm loc}\nolimits}(\phi_{v},0)$ for at least one finite place $v$.
@@
As explained in [17, p. 583], the results of [20] imply that if $\phi\in\Phi_{p}$, then the spaces $I^{\prime}_{\phi}$, resp. $J_{\phi}$ are non-trivial, proper subspaces of the domain, resp. of the codomain of $T^{\mathop{\rm loc}\nolimits}(\phi,0)$.
@@
[17, Thm. 2] Let $r$ be the restriction morphism of (36). Then, $r_{{\mathbb{C}}}\circ(\oplus_{\phi\in\Phi_{\mbox{\tiny{np}}}}\mathop{\rm Eis}\nolimits_{\phi}\oplus_{\phi\in\Phi_{\mbox{\tiny{p}}}}\mathop{\rm Eis}\nolimits^{\prime}_{\phi})$ induces an isomorphism
@@
    The case $\phi\in\Phi_{\mbox{\tiny{np}}}$ of the above theorem is proven in [17, pp. 581-583], and the case $\phi\in\Phi_{\mbox{\tiny{p}}}$ in [17, pp. 583-584].
@@
Let $k$ be a positive integer and let $\phi$ be a Hecke character of $F$ of infinity type $(k,-k-3)$. Suppose $\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}(\cdot)=|\cdot|_{{\mathbb{I}}_{\mathbb{Q}}}^{3}$.
@@
For any compact open $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is small enough, it endows $I^{K}_{\phi}$ with a pure Hodge structure, direct sum of 1-dimensional pure $\overline{{\mathbb{Q}}}$-Hodge structures
@@
[…]
In order to state the second consequence of Theorem 5.13 which will be of interest for us, denote, for $i\in\{0,\dots,3\}$
@@
For any compact open $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is small enough, it endows $I^{K}_{\theta\phi}$ with a pure Hodge structure, direct summand of the one on $H^{1}_{c,\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$. It is the direct sum of 1-dimensional pure $\overline{{\mathbb{Q}}}$-Hodge structures
@@
[…]
@@
Fix $\phi$ as in the above Corollary, and $K$ small enough. We will denote the pure Hodge structure having $I^{K}_{\phi}$ as underlying vector space by
@@
Suppose $\phi_{{\mathbb{Q}}}$ satisfies the same assumption as in Corollary 5.15, but $\phi$ is of infinity type $(-k-3,k)$. Then, using the local system $V^{\vee}_{k}$ instead of $V_{k}$, and employing Corollary 4.4, the same proof shows the existence of a sub-Hodge structure $\partial H^{2}(\phi)$ of degree-2 boundary cohomology, of Hodge type $(k+1,0)$ and of underlying vector space $I^{K}_{\phi}$.
@@ 6. A subspace of interior cohomology
The aim of this section is to construct specific sub-Hodge structures $H^{2}(\phi)_{!}$ of interior cohomology of Picard surfaces. These subspaces are associated with Hecke characters $\phi$ of the imaginary quadratic field $F$ satisfying the assumptions of our main theorem, Theorem 1, and provide indeed the last ingredient which will be needed to prove it in the next section.
@@ 6.1. The setup
[…]
We will consider those $\phi$’s such that the restriction $\phi_{{\mathbb{Q}}}$ of $\phi$ to ${\mathbb{Q}}$ satisfies
(45) $\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}$
@@
Let us discuss the existence of Hecke characters verifying the above assumption. Observe that an Hecke character $\phi$ of $F$ infinity type $(k,-k-3)$ can be written as
@@
where $\mathop{\rm IndUn}\nolimits$ denotes the unitary induction functor. Moreover, we have $\phi_{u,\infty}(z)=z^{-k-\frac{3}{2}}\overline{z}^{k+\frac{3}{2}}$ for $z\in{\mathbb{C}}^{\times}$.
Now, the proof of [4, Lemma 6.9.2] shows: $(a)$ that condition (45) is equivalent to asking that $\phi_{u}=\overline{\phi_{u}}^{-1}$; $(b)$ that such $\phi_{u}$’s verifying the latter condition do exist.
@@
(48) $L(\phi,-1)=0$
Under (45), the intertwining operator $T(\phi,s)$ may have a pole at $s=0$, but when (48) holds, this is not the case; hence, $T(\phi,0)$ is well defined.
@@
Let $k$ be a positive integer and $\phi$ be an Hecke character of $F$ of infinity type $(k,-k-3)$. If $\phi$ satisfies (45) and (48), then we denote the Langlands quotient of $I^{*}_{\phi}$ by
@@
    By the results of [20] (see also [29, 12.2]), asking that $\phi$ verifies (45) implies that $J^{*}_{\phi}$ is non-trivial.
@@
Observe that if we suppose that the sign $\epsilon:=\epsilon(-1)$ of the functional equation of $\phi$ satisfies
(49) $\epsilon=-1$
@@
is the sub-$\mathrm{G}({\mathbb{A}}_{f})$-module of interior cohomology $H_{!}^{\bullet}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})$ (cfr. (32)), defined as the direct sum of all the irreducible sub-$\mathrm{G}({\mathbb{A}}_{f})$-modules of $H^{\bullet}_{!}(S({\mathbb{C}}),V_{\lambda})$ isomorphic to the non-archimedean component of some cuspidal automorphic representation of $\mathrm{G}({\mathbb{A}})$. If we denote the set of these isomorphism classes by $\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}$, then by multiplicity one for the discrete spectrum of $\mathrm{G}({\mathbb{A}})$ ([29, Thm. 13.3.1]), cuspidal cohomology decomposes as
(50) $H^{\bullet}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})\simeq\bigoplus\limits_{\begin{subarray}{c}\pi=\pi_{f}\otimes\pi_{\infty}\\ \mbox{\tiny{s. t.}}\ \pi_{f}\in\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}\end{subarray}}\pi_{f}\otimes H^{\bullet}({\mathfrak{g}},\mathop{\tilde{\rm K}_{\infty}}\nolimits,\pi_{\infty}\otimes V_{\lambda,{\mathbb{C}}})$
In [18, p. 99], the following has been observed. It is a consequence of [29, Thms. 13.3.6, 13.3.7], as corrected in [31, Thm. 1.1], that under (49), there exists a $\mathrm{G}({\mathbb{A}}_{f})$-submodule $H^{2}(\pi(\phi)_{f})$ of $H^{2}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{k})$, whose local components are isomorphic to $J^{*}_{\phi,v}$ for infinitely many places $v$.
@@ 6.2. An application of Rogawski’s results
Given the Hecke character $\phi$, we are now going to construct the $\mathrm{G}({\mathbb{A}}_{f})$-submodule $H^{2}(\pi(\phi)_{f})$ mentioned above, by constructing a representation of $\mathrm{G}({\mathbb{A}})$ which we will show to be cuspidal and cohomological as an application of Rogawski’s results.
For this, we fix a choice of a Hecke character $\mu$ of $F$ whose restriction to ${\mathbb{Q}}$ is $\omega_{F|{\mathbb{Q}}}(\cdot)$. Moreover, we choose a couple $(\rho,\rho^{\prime})$ of characters on the norm-one subgroup $C_{F}$ of the idèle class group of $F$ as follows: define $\rho_{1}$ to be the restriction of $\mu\cdot\phi_{u}$ to $C_{F}$, and $\rho^{\prime}$ to be trivial. To such a couple, one attaches a character $\rho$ on $\mathrm{U}_{2}\times\mathrm{U}_{1}$ as in [31, p. 395].
@@
    is $v$ is inert or ramified in $F$, we define it as the Langlands quotient $J_{\phi,v}$ of $I_{\phi,v}$, (non-trivial under our hypothesis on $\phi$, by Rmk. 6.3);
@@
    if $v$ splits into places $w_{1}$ and $w_{2}$ in $F$, identifying $\mathrm{G}({\mathbb{Q}}_{v})\simeq\mathop{\rm GL}\nolimits_{3}(F_{w_{1}})\times F^{\times}_{w_{1}}$, and seeing $\phi_{v}$ as a character of $(F^{\times}_{w_{1}})^{4}$, we define it as the representation induced from the standard $(2,1)$ parabolic of $\mathop{\rm GL}\nolimits_{3}(F_{w_{1}})$ by the character
@@
For any finite place $v$, the unitarily induced representation called $\mathop{\rm ind}\nolimits(\eta_{v})$ in [31, p. 395] is nothing but our $I_{\phi_{v}}$, or more precisely, its restriction to a representation of $\mathrm{G}^{\nu=1}$. This follows by observing that in the notation of (47), the character called $\phi$ in loc. cit. becomes $\phi_{u}$. This shows that for $v$ inert or ramified, our notation $\pi^{n}(\rho_{v})$ is coherent with Rogawski’s one in [31, p. 395]. The same holds, by definition, in the case of split $v$ ([31, p. 396]).
@@
The discrete series representation $\pi(\phi)_{\infty}$ of $\mathrm{G}({\mathbb{R}})$ is defined as the representation $\pi^{s}(\rho_{\infty})$ denoted by $\pi^{s}(\rho)$ in [30, p. 77, first line].
@@
Let $\lambda=(k_{1},k_{2},c,r)$ be a character of $\mathrm{T}_{F}$. By [30, 3.2], the representation $\pi^{s}(\rho)$ is $V_{\lambda}$-cohomological in degree 2 when either $k_{2}=0$ or $k_{1}=k_{2}$. In the first case, it coincides with the discrete series representation denoted by $\pi^{+}$ in loc. cit.; in the second case, it coincides with the one denoted by $\pi^{-}$.
To see how this follows from loc. cit., observe that in our notation, the restriction to $\mathrm{G}^{\nu=1}$ of the representation $V_{\lambda}$ is identified with a triple $(k_{1},k_{2},c)$, whereas in [30, p. 79] it is identified with a triple $(m,r,n)$. Now, under the change of variables between the two parametrizations, one sees that $k_{2}=0$ if and only if $r-n=1$, and that $k_{1}=k_{2}$ if and only if $m-r=1$.
@@
Let $k$ be a positive integer and $\phi$ an Hecke character of $F$ of type $(k,-k-3)$, satisfying
@@
    The automorphic representation $\pi(\phi)$ belongs to the $A$-packet $\Pi(\rho)$ of [31, p. 396], and by construction, the number $n(\pi)$ of loc. cit. of places $v$ where $\pi(\phi)_{v}$ is isomorphic to $\pi^{s}(\rho_{v})$ is equal to 1 (since this happens only for $v=\infty$). Now write $\phi_{R}$ for the character denoted $\phi$ in loc. cit.. By Remark 6.5, we have $\phi_{R}=\phi^{(1)}_{u}$, so that by (46)
    $L\left(\phi_{R},\frac{1}{2}\right)=L\left(\phi\cdot|\cdot|_{L}^{-\frac{3}{2}},\frac{1}{2}\right)=L\left(\phi,-1\right)$
    which shows that our hypothesis $\epsilon(\phi)=-1$ is precisely equivalent to $\epsilon(\frac{1}{2},\phi_{R})=-1$. By [31, Thm. 1.1], we obtain that the multiplicity of $\pi(\phi)$ in the discrete spectrum of $\mathrm{G}$ is 1. By [29, Thm. 13.3.6 (a)], it is actually a representation appearing in the cuspidal spectrum.
@@
    where each space $\mathop{\rm Hom}\nolimits_{\mathrm{G}({\mathbb{A}}_{f})}(\pi_{f},H^{\bullet}_{!}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}))$ is endowed with a canonical $\overline{{\mathbb{Q}}}$-Hodge structure. Now consider the submodule corresponding to $\pi=\pi(\phi)$. By Remark 6.7, $\pi(\phi)_{\infty}$ is $V_{k}$-cohomological in degree 2. By the comparison isomorphism between singular interior cohomology and étale interior cohomology, the dimension of the Hodge structure

### v4

@@ 5.3. Automorphic description of Eisenstein cohomology
To describe the main result of this section, we consider a modification of the intertwining operators $T(\phi,s)$ introduced in the previous subsection.
@@
Let $\lambda$ be a character of $\mathrm{T}^{\prime}$ and $w$ be an element of the Weyl group $W$ of $G_{F}$ of length $\geq 2$. Let $\phi$ be a Hecke character of $F$ of type $w\star\lambda$. Let $s\in{\mathbb{C}}$ belong to the half-plane of convergence of the Euler product of $L(\phi,s)$. For every place $v$, we define
@@
    The definition makes sense because if $s$ satisfies the hypotheses, i.e. if $\mbox{Re}\ s$ satisfies the bounds of Lemma 5.4(1), then both the operator $T(\phi_{v},s)$ and the quantity $c_{v}(\phi,s)$ are indeed defined and non-zero (cfr. [23, Lem. 2.3.1]).
@@
For the remainder of this section, we assume that $\lambda$, $w$, and $\phi$ are as in the preceding definition.
@@
    We denote $I^{\prime}_{\phi}$ the submodule of $I_{\phi}$ generated by those tensors, whose component at $v$ belongs to the kernel of $T^{\mathop{\rm loc}\nolimits}(\phi_{v},0)$ for at least one finite place $v$. We denote $I^{\prime,*}_{\phi}:=I^{\prime}_{\phi}\otimes I_{\phi_{\infty}}$.
@@
As explained in [23, p. 583], the results of [26] imply that if $\phi\in\Phi_{p}$, then the spaces $I^{\prime}_{\phi}$, resp. $J_{\phi}$ are non-trivial, proper subspaces of the domain, resp. of the codomain of $T^{\mathop{\rm loc}\nolimits}(\phi,0)$.
Definition 5.12.
  1. (1)
[…]
  2. (2)
[…]
@@
[23, Thm. 2] Let $r$ be the restriction morphism of (41). Then, $r_{{\mathbb{C}}}\circ(\oplus_{\phi\in\Phi_{\mbox{\tiny{np}}}}\mathop{\rm Eis}\nolimits_{\phi}\oplus_{\phi\in\Phi_{\mbox{\tiny{p}}}}\mathop{\rm Eis}\nolimits^{\prime}_{\phi})$ induces an isomorphism
@@
    The case $\phi\in\Phi_{\mbox{\tiny{np}}}$ of the above theorem is proven in [23, pp. 581-583], and the case $\phi\in\Phi_{\mbox{\tiny{p}}}$ in [23, pp. 583-584].
@@
Let $k$ be a positive integer and let $\phi$ be a Hecke character of $F$ of infinity type (k,k3)(k,-k-3). Suppose $\phi_{{\mathbb{Q}}}\cdot\omega_{F|{\mathbb{Q}}}(\cdot)=|\cdot|_{{\mathbb{I}}_{\mathbb{Q}}}^{3}$.
@@
For any compact open $K\subset\mathrm{G}({\mathbb{A}}_{f})$ which is small enough, it endows $I^{K}_{\phi}$ with a pure Hodge structure, isomorphic to a direct sum of 1-dimensional pure $\overline{{\mathbb{Q}}}$-Hodge structures
@@
Let $\phi$ be as in the hypotheses and observe that by putting $\lambda=(k,0)$, its type can be written as $\mbox{(1\ 2\ 3\ )}\star\lambda$. Then, by Corollary 5.6, $c(\phi,s)$ has no pole in $s=0$. Hence, by Theorem 5.13, $I_{\phi}$ appears as a submodule of degree 2 Eisenstein cohomology. Once one remembers that by Corollary 4.3, the pure 1-dimensional $F$-Hodge structure $I_{k,-k-3}$ has Hodge type $(1,k+2)$, the description of the Hodge structure on $I^{K}_{\phi}$ comes from the isomorphism (39). ∎
To state the second consequence of Theorem 5.13 which will be of interest to us, denote, for $i\in\{0,\dots,3\}$
@@
For any compact open $K\subset\mathrm{G}({\mathbb{A}}_{f})$ that is small enough, it endows $I^{K}_{\theta\phi}$ with a pure Hodge structure, a direct summand of the one on $H^{1}_{c,\mathop{\rm Eis}\nolimits}(\mathrm{S}_{K}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}})$. It is the direct sum of 1-dimensional pure $\overline{{\mathbb{Q}}}$-Hodge structures
@@
Note that if $\phi$ is of infinity type (k,k3)(k,-k-3), that is, of type $\mbox{(1\ 2\ 3\ )}\star(k,0)$, then $\theta\phi$ is of type $\mbox{(1\ 2\ )}\star(k,0)$, that is, of infinity type (k+1,k2)(k+1,-k-2). Then, the statement is proven by applying Corollary 4.8 and Theorem 5.13, reasoning analogously to the proof of Corollary 5.15, and noticing that Corollary 4.8 exhibits $I_{\theta\phi}$ as a direct summand of boundary cohomology. ∎
@@
Fix $\phi$ as in the above corollary and $K$ small enough. We will denote the pure Hodge structure having $I^{K}_{\phi}$ as underlying vector space by
@@
Suppose $\phi_{{\mathbb{Q}}}$ satisfies the same assumptions as in Corollary 5.15, but $\phi$ is of infinity type (k3,k)(-k-3,k). Then, using the local system $V^{\vee}_{k}$ instead of $V_{k}$, and employing Corollary 4.4, the same proof shows the existence of a sub-Hodge structure $\partial H^{2}(\phi)$ of degree-2 boundary cohomology, of Hodge type $(k+1,0)$ and of underlying vector space $I^{K}_{\phi}$.
@@ 6. A subspace of interior cohomology
The aim of this section is to construct specific sub-Hodge structures $H^{2}(\phi)_{!}$ of interior cohomology of Picard surfaces. These subspaces are associated with Hecke characters $\phi$ of the imaginary quadratic field $F$ satisfying the assumptions of our Theorem 1. They indeed provide the last ingredient, which will be needed in the next section to prove the theorem.
@@ 6.1. The setup
We fix a positive integer $k$. We will consider Hecke characters $\phi$ of $F$ of infinity type (k,k3)(k,-k-3). In particular, the weight of such a $\phi$ is 3-3 and the center of its functional equation is s=1s=-1.
For any such $\phi$, we identify it with a Hecke character on the maximal torus $\mathrm{T}^{\prime}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$ according to the conventions of Subsection 3.1. Moreover, we fix any extension of $\phi$ to a Hecke character of the maximal torus $\mathrm{T}$ of $\mathrm{G}$, and denote it by the same symbol; the context will make it clear when we need to see $\phi$ as a Hecke character of $\mathrm{T}$ instead of just $F$.
The first hypothesis on our $\phi$’s will be that the restriction $\phi_{{\mathbb{Q}}}$ of $\phi$ to $\mathbb{Q}$} satisfies
(50) $\phi_{{\mathbb{Q}}}=\omega_{F|{\mathbb{Q}}}|\cdot|^{3}_{{\mathbb{I}}_{{\mathbb{Q}}}}.$
@@
Let us discuss the existence of Hecke characters verifying the above assumption. Observe that using the element $\delta$ defined in (15), a Hecke character $\phi$ of $F$ of infinity type (k,k3)(k,-k-3) can be written as
@@
where $\mathop{\rm IndUn}\nolimits$ denotes the unitary induction functor. Moreover, we have
$\phi_{u,\infty}(z)=z^{-k-\frac{3}{2}}\overline{z}^{k+\frac{3}{2}}\quad\text{for}\quad z\in{\mathbb{C}}^{\times}.$
Now, define $\phi^{\perp}$ and $\phi_{u}^{\perp}$ as in (28). Then, the proof of [7, Lem. 6.9.2] shows that:
  1. (a)
    condition (50) is equivalent to asking that $\phi_{u}=\phi^{\perp}_{u}$, hence to asking
    (53) $\phi^{\perp}=\phi|\cdot|^{-3}_{{\mathbb{I}}_{F}},$
  2. (b)
    such $\phi_{u}$’s verifying $\phi^{\perp}_{u}=\phi_{u}$ do exist.
@@
(54) $L(\phi,-1)=0.$
Lemma 5.3 tells us that under (50), the intertwining operator $T(\phi,s)$ of Definition 5.1 may have a pole at $s=0$, but the same Lemma also implies that when (54) holds, no pole can appear. Hence, under this hypothesis, $T(\phi,0)$ is well defined.
@@
Let $k$ be a positive integer and $\phi$ be a Hecke character of $F$ of infinity type (k,k3)(k,-k-3). If $\phi$ satisfies (50) and (54), then we denote
@@
    Let $p$ be a prime that is inert or ramified in $F$. Then $J_{\phi,p}$ is the Langlands quotient of $I_{\phi,p}$, that is, its unique irreducible quotient as a $\mathrm{G}({\mathbb{Q}}_{p})$-module. By the results of [26] (as explained in [36, Sect. 12.2]), the module $J_{\phi,p}$ is a proper quotient of $I_{\phi,p}$ precisely under the hypothesis (50).
@@
Now suppose that $\phi$ satisfies (50), or equivalently (53). By Remark 3.2, if the sign $\epsilon(\phi)$ satisfies
(55) $\epsilon(\phi)=-1$
@@
is the $\mathrm{G}({\mathbb{A}}_{f})$-submodule of interior cohomology $H_{!}^{\bullet}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})$ (cfr. (36)), defined as the direct sum of all those irreducible $\mathrm{G}({\mathbb{A}}_{f})$-submodules of $H^{\bullet}_{!}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})$ which are isomorphic to the non-archimedean component of some cuspidal automorphic representation of $\mathrm{G}({\mathbb{A}})$. If we denote the set of these isomorphism classes by $\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}$, then by multiplicity one for the discrete spectrum of $\mathrm{G}({\mathbb{A}})$ ([36, Thm. 13.3.1]), cuspidal cohomology decomposes as
(56) $H^{\bullet}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{\lambda,{\mathbb{C}}})\simeq\bigoplus\limits_{\begin{subarray}{c}\pi=\pi_{f}\otimes\pi_{\infty}\\ \mbox{\tiny{s. t.}}\ \pi_{f}\in\mathcal{C}_{\mathop{\rm\tiny{cusp}}\nolimits}\end{subarray}}\pi_{f}\otimes H^{\bullet}({\mathfrak{g}},\mathop{\tilde{\rm K}_{\infty}}\nolimits,\pi_{\infty}\otimes V_{\lambda,{\mathbb{C}}}).$
In [24, p. 99], the following observation is made. As a consequence of [36, Thms. 13.3.6, 13.3.7], corrected in [38, Thm. 1.1], it follows that under assumption (55), there exists a $\mathrm{G}({\mathbb{A}}_{f})$-submodule
(57) $H^{2}(\pi(\phi)_{f})\quad\text{of}\quad H^{2}_{\mathop{\rm\tiny{cusp}}\nolimits}(S({\mathbb{C}}),V_{k}),$
whose local components are isomorphic to $J_{\phi,v}$ for infinitely many places $v$. We explain this in the following subsection.
@@ 6.2. An application of Rogawski’s results
Given the Hecke character $\phi$, we are now going to construct the $\mathrm{G}({\mathbb{A}}_{f})$-submodule $H^{2}(\pi(\phi)_{f})$ mentioned in equation (57). This will amount to constructing a representation of $\mathrm{G}({\mathbb{A}})$ which we will show to be cuspidal and cohomological by applying Rogawski’s results.
For this, we fix a choice of a Hecke character $\mu$ of $F$ whose restriction to $\mathbb{Q}$} is $\omega_{F|{\mathbb{Q}}}(\cdot)$. Moreover, we choose a couple of characters $(\rho,\rho^{\prime})$ on the norm-one subgroup $C_{F}$ of the idèle class group of $F$ as follows: define $\rho_{1}$ to be the restriction of $\mu\cdot\phi_{u}$ to $C_{F}$, and $\rho^{\prime}$ to be trivial. To such a couple, one attaches a character $\rho$ on $\mathrm{U}_{2}\times\mathrm{U}_{1}$ as in [38, p. 395].
@@
    is $v$ is inert or ramified in $F$, then we define $\pi^{n}(\rho_{v})$ as the Langlands quotient $J_{\phi,v}$ of $I_{\phi,v}$, (see Remark 6.3);
@@
    if $v$ splits into places $w_{1}$ and $w_{2}$ in $F$, then identifying $\mathrm{G}({\mathbb{Q}}_{v})\simeq\mathop{\rm GL}\nolimits_{3}(F_{w_{1}})\times F^{\times}_{w_{1}}$, and seeing $\phi_{v}$ as a character of $(F^{\times}_{w_{1}})^{4}$, we define $\pi^{n}(\rho_{v})$ as the representation induced from the standard $(2,1)$ parabolic of $\mathop{\rm GL}\nolimits_{3}(F_{w_{1}})$ by the character
@@
For any finite place $v$, the unitarily induced representation called $\mathop{\rm ind}\nolimits(\eta_{v})$ in [38, p. 395] is nothing but our $I_{\phi_{v}}$, or more precisely, its restriction to a representation of $\mathrm{G}^{\nu=1}$. This follows by observing that in the notation of (52), the character called $\phi$ in loc. cit. becomes the character $\phi_{u}$ defined in (51). This shows that for $v$ inert or ramified, our notation $\pi^{n}(\rho_{v})$ is coherent with Rogawski’s one in [38, p. 395]. The same holds, by definition, in the case of split $v$ ([38, p. 396]).
@@
The discrete series representation $\pi(\phi)_{\infty}$ of $\mathrm{G}({\mathbb{R}})$ is defined as the representation $\pi^{s}(\rho_{\infty})$ denoted by $\pi^{s}(\rho)$ in [37, p. 77, first line].
@@
Let $\lambda=(k_{1},k_{2},c,r)$ be a character of $\mathrm{T}_{F}$. By [37, 3.2], the representation $\pi^{s}(\rho)$ is $V_{\lambda}$-cohomological in degree 2 when either $k_{2}=0$ or $k_{1}=k_{2}$. In the first case, it coincides with the discrete series representation denoted by $\pi^{+}$ in loc. cit.; in the second case, it coincides with the one denoted by $\pi^{-}$.
To see how this follows from loc. cit., observe that in our notation, the restriction to $\mathrm{G}^{\nu=1}$ of the representation $V_{\lambda}$ is identified with a triple $(k_{1},k_{2},c)$, whereas in [37, p. 79] it is identified with a triple $(m,r,n)$. Now, under the change of variables between the two parametrizations, one sees that $k_{2}=0$ if and only if rn=1r-n=1, and that $k_{1}=k_{2}$ if and only if mr=1m-r=1.
For the following statement, recall the level $K$ Hecke algebra defined in Remark 4.5. Moreover, recall that if $\phi$ satisfies (50), or equivalently (53), then by Remark 3.2 there is a well-defined sign $\epsilon(\phi)$.
@@
Let $k$ be a positive integer and $\phi$ a Hecke character of $F$ of type (k,k3)(k,-k-3), satisfying
@@
    The automorphic representation $\pi(\phi)$ belongs to the $A$-packet $\Pi(\rho)$ of [38, p. 396], and by construction, the number $n(\pi)$ of loc. cit. of places $v$ where $\pi(\phi)_{v}$ is isomorphic to $\pi^{s}(\rho_{v})$ is equal to 1 (since this happens only for $v=\infty$). Now write $\phi_{R}$ for the character denoted $\phi$ in loc. cit.. By Remark 6.5, we have $\phi_{R}=\phi^{(1)}_{u}$, so that by (51)
    $L\left(\phi_{R},\frac{1}{2}\right)=L\left(\phi\cdot|\cdot|_{{\mathbb{I}}_{F}}^{-\frac{3}{2}},\frac{1}{2}\right)=L\left(\phi,-1\right)$
    which shows that our hypothesis $\epsilon(\phi)=-1$ is precisely equivalent to $\epsilon(\frac{1}{2},\phi_{R})=-1$. By [38, Thm. 1.1], we obtain that the multiplicity of $\pi(\phi)$ in the discrete spectrum of $\mathrm{G}$ is 1. By [36, Thm. 13.3.6 (a)], it is actually a representation appearing in the cuspidal spectrum.
@@
    where each space $\mathop{\rm Hom}\nolimits_{\mathrm{G}({\mathbb{A}}_{f})}(\pi_{f},H^{\bullet}_{!}(\mathrm{S}({\mathbb{C}}),V_{k,\overline{{\mathbb{Q}}}}))$ is endowed with a canonical $\overline{{\mathbb{Q}}}$-Hodge structure. Now consider the submodule corresponding to $\pi=\pi(\phi)$. By Remark 6.7, $\pi(\phi)_{\infty}$ is $V_{k}$-cohomological in degree 2. By the comparison isomorphism between singular interior cohomology and étale interior cohomology (see e.g. [10, Sect. 1.7], the dimension of the Hodge structure